Accounting please help me with this question ASAP

(Ratio analysis over time)
The following information comes from the accounting records of Hercep Ltd. for the first three years of its existence:



2018
2019
2020
Statement of Financial Position





Assets





   Cash
$22,300
$19,200
$25,100
   Accounts receivable
67,400
58,500
149,500
   Inventory
113,000
139,000
224,000
   Capital assets (net)
425,000
455,000
468,000
   Other assets
235,000
205,000
265,200


$862,700
$876,700
$1,131,800
Liabilities and equity





   Accounts payable
$117,000
$58,500
$117,000
   Long-term debt
184,000
234,000
468,000
   Common shares
517,000
517,000
517,000
   Retained earnings
44,700
67,200
29,800


$862,700
$876,700
$1,131,800
Statement of Earnings





   Sales
$700,000
$815,000
$915,000
   Cost of goods sold
(434,000)
(554,200)
(640,500)
   Other expenses
(160,000)
(235,000)
(233,000)


106,000
25,800
41,500
   Income tax
(31,800)
(7,740)
(12,450)
   Net income
$74,200
$18,060
$29,050

Based on above information, analyze the changes in the company’s profitability and liquidity, in addition to the management of accounts receivable and inventory from 2018 to 2020. (Round answers to 1 decimal place, e.g. 13.5% or 13.5.)



2018
2019
2020
Sales
  %
  %
  %
Cost of goods sold
  %
  %
  %
Gross margin
  %
  %
  %
Other expenses
  %
  %
  %
Income taxes
  %
  %
  %
Net income
  %
  %
  %




2018
2019
2020
Current Ratio
: 1
: 1
: 1
Quick Ratio
: 1
: 1
: 1
A/R Turnover
times
times
times
Average Collection Period
days
days
days
Inventory Turnover
times
times
times
Days to Sell Inventory
days
days
days


    

Based on the above information, analyze the company’s use of leverage from 2018 to 2020. (Round answers to 1 decimal place, e.g. 15.1%.)



2018
2019
2020
Debt to Equity
  %
  %
  %
Return on Assets
  %
  %
  %
Return on Equity
  %
  %
  %
Answer #1

SOLUTIONS

2018 2019 2020
Sales 100   % 100   % 100   %
Cost of goods sold 62   % 68   % 70   %
Gross margin 38   % 32   % 30   %
Other expenses 23   % 29   % 25   %
Income taxes 4.5   % 0.94   % 1.36   %
Net income 10.6   % 2.21   % 3.17   %

COST OF GOODS SOLD % = COST OF GOODS SOLD / SALES

2018 = $434000/$700000 =62%

2019 = $554200/$815000 = 68%

2020 = $640500/$915000 = 70%

INCOME TAX %

INCOME TAX / SALES = INCOME TAX %

2018 = $31800/$700000 =4.5%

2019 = $7740/$815000=0.94%

2020 = $12450/$915000= 1.36%

GROSS MARGIN

REVENUE - COST OF GOODS SOLD/ REVENUE

2018 = $700000-$434000/$700000 = 38%

2019 = $815000-$554200/$815000 = 32%

2020 = $915000-640500/915000 = 30%

A company's gross profit margin percentage is calculated by first subtracting the cost of goods sold (COGS) from the net sales (gross revenues minus returns, allowances, and discounts). This figure is then divided by net sales, to calculate the gross profit margin in percentage terms

OTHER EXPENSES %

OTHER EXPENSES / SALES = OTHER EXPENSES %

2018 = $160000/$700000 =23%

2019 = $235000/$815000= 29%

2020 = $233000/$915000= 25%

NET INCOME%

NET INCOME / SALES = NET INCOME %

2018 = 74200/700000= 10.6

2019 = 18060/815000= 2.21%

2020 = 29050/915000= 3.17%

2018 2019 2020

1) Current Ratio 1.73 3.70 3.40   

2) Quick Ratio 0.762 1.328 1.49

3) A/R Turnover 10.38 12.94 8.79

4) AVERAGE COLLECTION PERIOD 35.16 28.20 41.52

5) INVENTORY TURN OVER 3.840 4.39 3.528

6) DAYS TO SELL INVENTORY 95.05 83.14 103.45

1) CURRENT RATIO = CURRENT ASSET/ CURRENT LIABILITY

CURRENT ASSET = CASH +A/R +INVENTORY

CURRENT LIABILITY= A/P

2) QUICK RATIO =CURRENT ASSET-INVENTORY/CURRENT LIABLITY

3) A/R Turnover= Net Annual Credit Sales ÷ ((Beginning Accounts Receivable + Ending Accounts Receivable) / 2)

2018 = 700000/67400 =10.38 TIMES

2019=815000/58500+67400/2 =12.94TIMES

2020= 915000/149500+58500/2 = 8.79 TIMES

4) AVERAGE COLLECTION PERIOD

365DAYS/A/R TURNOVER

2018 =365/10.38TIMES =35.16

2019= 365/ 12.94 TIMES= 28.20

2020= 365/ 8.79 TIMES= 41.52

5) INVENTORY TURN OVER

COST OF GOODS SOLD / AVERAGE INVENTORY

6) DAYS TO SELL INVENTORY

DSI= 1 / inventory turnover​×365 days

Debt to equity ratio   =   Total liabilities / Total stockholder's equity * 100
2018 $ 301000/ $561700 * 100 53.58%
2019 $ 292500/ $584200 * 100 50.06%
2020 $585000 / $ 546800* 100 106.98%
*Calculations for Total liabilities:
2018 2019 2020
Accounts payable $117000 $58,500 $117,000
Long term debt $184,000 $234,000 $468000
Total liabilities $301000 $292500 $585000
*Calculation of Total stockholder's equity:
2018 2019 2020
Common shares $517000 $517000 $517000
Retained earnings $44700 $67200 $29800
Total stockholder's equity $561700 $584200 $546800
Ans. Return on assets =   Net income / Total assets * 100
2018 $74200 / $862700 * 100 8.60%
2019 $18060 / $876700 * 100 2.06%
2020 $29050/ $1131800 * 100 2.566%
Ans. Return on equity =   Net income / Total stockholder's equity * 100
2018 $ 74200/ $561700 * 100 13.2%
2019 $18060 / $584200 * 100 3.09%
2020 $29050 / $546800 * 100 5.312%